1.  A charge of Q1 of 2 x 10-6 C is placed on a line between two charges Q2 of 1 x 10-6 C and Q3 of -6 x 10-6 C. Q1 is 10 x 10-4 m from Q2 and 20 x 10-4 m from Q3. What is the net force on Q1?


1
Expert's answer
2022-03-21T12:46:49-0400

Force

"F_3=\frac{kq_1q_3}{r^2}"


"F_3=-\frac{9\times10^{9}\times2\times10^{-6}\times6\times10^{-6}}{(2\times10^{-3})^2}=-27000N"

"F_2=\frac{kq_1q_2}{r^2}"


"F_1=-\frac{9\times10^{9}\times2\times10^{-6}\times1\times10^{-6}}{(1\times10^{-3})^2}=18000N"

"F_{net}=F_1+F_2"

"F_{net}=18000-27000=-9000N"